Pittman Company is a small but growing manufacturer of telecommunications equipment. The company has no sales force of its owI say its just plain robbery, retorted Karl. And I also say its time we dumped those guys and got our own sales force. C4. Compute the degree of operating leverage that the company would expect to have at the end of next year assuming: a. The ag
Answer #1
Solution 1:
15% commission 20% commission own sales force
Sales 23500000 23500000 23500000
Variable expenses:
Manufacturing 10575000 10575000 10575000
Commission 3525000 4700000 1762500
Total variable expenses 14100000 15275000 12337500
Contribution margin 9400000 8225000 11162500
Variable expenses ratio 60.00% 65.00% 52.50%
Contribution margin ratio 40.000% 35.000% 47.500%
Fixed expenses:
Manufacturing overhead 3290000 3290000 3290000
Marketing 164500 164500 3689500
Administrative 2100000 2100000 1991900
Interest 822500 822500 822500
Total fixed expenses 6377000 6377000 9793900
Income before income taxes 3023000 1848000 1368600
Income taxes(30%) 906900 554400 410580
Net income 2116100 1293600 958020
Break Even Point in dollars (Total Fixed Expenses/Controbution margin ratio) 15942500 18220000 20618737
Solution 2:
Income before tax as per income statement 3023000
Total Fixed Expenses (when 20% commission) 6377000
Total 9400000
/ Contribution margin ratio (When 20% commission) 35.000%
Sales required to generate target profit 26857143
Solution 3:
Let Sales Volume be "X"
Total Cost when commission 20% = Total Cost for Own sales Force
0.6500*X +6377000 = 0.5250*X +9793900
0.125*X = 3416900
X = 3416900/0.125
X = 27335200
Hence, Sales volume required is $27,335,200.
Solution 4:
15% commission 20% commission own sales force
Contribution margin 9400000 8225000 11162500
Income before taxes 3023000 1848000 1368600
Degree of Operating Leverage (Contribution margin/Income before taxes) 3.11 4.45 8.16